how do you make 3/5 into a decimal rounded to the nearest thousandth and into a percent rounded to the nearest tenth of a percent
Answer by Theo(13342)

how do you make 3/5 into a decimal rounded to the nearest thousandth and into a percent rounded to the nearest tenth of a percent
-----
you divide 3 by 5 to get .6
-----
.6 doesn't need rounding to the nearest thousandth because it is an exact value when it is to the nearest tenth.
-----
if you rounded to the nearest integer, it would be 1.0
-----
if you must show it to the nearest thousandths, then you can show it as .600 because 1/1000 = .001 which is 3 decimal places.
-----
to make .6 into a percent you multiply it by 100%
.6 * 100% = 60%
-----
if you had a number like 7/11 then that would be a different story.
to convert 7/11 into a decimal you divide 7 by 11.
your answer would be .636363636363636..... and on and on and on.
now if you wish to round this to the nearest thousandths you would want to round it to 3 decimal places.
in order to do that you would take 4 decimal places and examine the 4th decimal place.
your number is .6363
the 4th decimal place is 3 which is less than 5 so you truncate to the 3d decimal place to get .636
-----
on the other hand, if your number was 15/7 and you wanted to convert it into a decimal, then you would divide 15/7 to get 2.142857143...
now if you want to round this to the nearest thousandths, you would want to round it to 3 decimal places because 1/1000 = .001
in order to do that you would take the first 4 decimal places and examine them.
your number would be 2.1428 to the 4th decimal place.
since the 4th decimal place is 8 which is greater than or equal to 5, you would add 1 to the 3d decimal place to get 2.143
-----
2.143 expressed as a percent would be 2.143 * 100% = 214.3%
-----